# Low Mass Naked Singularities from Dark Core Collapse

### v1

@@ 1 Introduction
Recent gravitational wave events GW190425 [1] and GW190814 [2] have suggested the existence of near-solar-mass collapsed objects which cannot be formed via stellar evolution. These two GW events have also opened up the possibility of reliable future detections of near-solar mass collapsed objects with the advanced Laser Interferometer Gravitational-Wave Observatory (aLIGO; [3]). An important question of physics and astronomy is how such a low mass collapsed object could form. One possibility is such near-solar mass objects are primordial black holes (PBHs; [4]), although there are observational constraints on PBHs [5]. An alternative is the capture of a PBH by a stellar host could transmute the host into a low mass black hole [6]. As the probability of such a capture is small [7], one needs to look for further alternatives. According to one such scenario, dark matter (DM) particles could accumulate around the center of a neutron star and form a mini BH by collapse, which consumes the whole star by accreting matter from it [8]. Thus, the whole neutron star is transmuted into a low mass BH.
@@
In general, a collapsed object is considered to be a BH due to the cosmic censorship conjecture (CCC) proposed by Penrose [9]. According to the CCC, the nature abhors the singularity. Otherwise, one can see the singularity from infinity, in principle, and such a collapsed object is called as a naked singularity. It was shown [10] that the initial condition of gravitational collapse determines the nature of the end product (BH or naked singularity) after completion of the gravitational collapse. For example, if the collapse is initiated by a completely homogeneous pressureless dust, with the density being constant from the center to the boundary of the collapsing cloud, the singularity will be hidden inside the event horizon, and the end product will be a BH. On the other hand, if the gravitational collapse is induced by an inhomogeneous dust, the end product can be visible as a naked singularity. Moreover, for an inhomogeneous collapsing cloud with a non-zero pressure, a naked singularity can arise [11]. Note that, as a perfectly homogeneous collapse is an idealistic scenario (e.g., stellar density changes in the radial direction), a gravitational collapse of a star may give rise to a naked singularity as the end product. In fact, Sgr A* is recently proposed as a 1st type of Joshi-Malafarina-Narayan (JMN-1 [10]) naked singularity [12]. Earlier, it was shown that M87* could be a Kerr-Taub-NUT naked singularity [13] or a Kerr naked singularity with the value of Kerr parameter within $4.5$ to $6.5$ [13]. Similarly, GRO J1655-40 could be a Kerr-Taub-NUT naked singularity [14]. In fact, it is theoretically shown in several papers [15, 11, 17, 18, 20, 19, 16] that the naked singularities are viable end products of collapse. In a recent paper [21], it is also indicated that many of the proposed primordial BHs could actually be primordial naked singularities. There are some effects already predicted [23, 22, 11] to distinguish a naked singularity from a BH, by which it may be possible to identify near- and sub-solar mass naked singularities in future.
@@
The nature of DM is not yet known. The interaction between the DM particles and the regular baryonic matter is via gravity, although there could be an extremely small additional interaction (with a very small cross section). There are several proposed DM candidates and some of them could be asymmetric or non-self-annihilating (see Fig. 1 of [24] for details). Let us consider that the non-self-annihilating DM particles enter a cosmic object/host (e.g., neutron star, white dwarf, etc.). They lose energy due to the scattering with the baryonic matter of the host. For example, in case of the neutron star, the DM particles lose energy due to the scattering with the neutron. Now, the captured DM particles are thermally distributed in a very small region of the core of the host due to the large baryon density [25]. Joining of the DM particles in an isothermal sphere around the core is known as thermalization [26]. For neutron stars, the DM reaches thermal equilibrium with the neutrons, and the captured DM particles quickly occupy a very small region near the core of the host [^4] We refer to Secs. III and IV of [25] for more details.. If a sufficiently large number of DM particles accumulate at the core of the host, a tiny collapsed object may form by the gravitational collapse [27]. In all the papers (e.g., [8, 27, 26, 25], and references therein) related to the dark core collapse, to the best of our knowledge, it has been presumed that the tiny collapsed object should be a BH, which may not be true. In this paper, we propose that depending on the initial condition (or, the nature) of the DM particles cloud, the tiny collapsed object could be either a BH or a naked singularity. For example, if the accumulated DM cloud is inhomogeneous in nature with a non-zero, tangential pressure, an object-like endoparasitic naked singularity [^5] Although this naked singularity is primarily hidden inside a stellar host, we call it a naked singularity in this sense that the singularity is not hidden inside the event horizon. For an object-like naked singularity, the event horizon does not form at all. could form at the core of the stellar host. The endoparasitic naked singularity should start accreting matter from its host immediately after its formation, and transmute the entire host into a naked singularity. Note that it is recently proposed that the gravitational collapse also gives rise to the primordial naked singularities [21] instead of primordial BHs depending on the nature of the collapsing matter. If such a primordial naked singularity is captured by a stellar object, it may transmute into a collapsed object. However, here we discuss only the transmutation led by an endoparasitic naked singularity formed by the accumulation of DM particles.
@@
This paper is organized as follows. We discuss the Chandrasekhar limit in Sec. 2. Sec. 3 describes the formation of an endoparasitic collapsed object inside a neutron star. We show in Sec. 4 that the endoparasitic collapsed object could be a BH or a naked singularity depending on the initial conditions of the DM cloud. In Sec. 5 we discuss the transmutation timescale, and we argue on the nature (BH or naked singularity) of the transmuted collapsed object in Sec. 6. We constrain a parameter space using the ‘missing pulsar’ problem in Sec. 7, and conclude in Sec. 8.
@@ 2 Chandrasekhar limit
$E_{\rm F}\sim\hbar cn^{1/3}\sim\hbar c\frac{N^{1/3}}{R}$ (1)
@@
$E=E_{\rm F}+E_{\rm G}=\hbar c\frac{N^{1/3}}{R}-\frac{GNm^{2}}{R}.$ (3)
@@
$N_{\rm Ch}^{\rm f}\sim\left(\frac{\hbar c}{Gm^{2}}\right)^{3/2}=\left(\frac{M_{\rm P}}{m}\right)^{3}=1.8\times 10^{51}\left(\frac{100~{}{\rm GeV}}{m}\right)^{3},$ (4)
@@
$E=\hbar c\frac{1}{R}-\frac{GNm^{2}}{R}.$ (5)
@@
$N_{\rm Ch}^{\rm b}\sim\left(\frac{\hbar c}{Gm^{2}}\right)=\left(\frac{M_{\rm P}}{m}\right)^{2}=1.5\times 10^{34}\left(\frac{100~{}{\rm GeV}}{m}\right)^{2}.$ (6)
For both fermions and bosons, the gravitational collapse eventually leads to form a collapsed object. Comparing Eqs. (4) and (6), one can see that bosons usually collapse more easily than fermions.
@@ 3 Formation of Endoparasitic Collapsed Objects inside Neutron stars
$C=\sqrt{\frac{6}{\pi}}\frac{\rho_{\rm DM}}{0.45m_{\chi}}{2GM_{\rm h}R_{\rm h}\over\bar{v}_{\chi}}.\left[1-\exp{\left(-\frac{\sigma}{\sigma_{\rm cr}}\right)}\right]\left[1-\frac{1-\exp{(-\xi)}}{\xi}\right]$ (7)
@@
where $N$ is the total number of DM particles accumulated inside the neutron star. $N_{\rm Ch}$ (see Eqs. 4 and 6) and $N_{\rm self}$ are the Chandrasekhar limit and the number of DM particles required for initiating the self-gravitating collapse, respectively [25]. If the total number ($N$) of fermions(f)/bosons(b) increases beyond $N^{\rm f}_{\rm Ch}$ or $N^{\rm b}_{\rm Ch}$ [25], the gravitational energy dominates over the total particle energy, and, the gravitational collapse occurs to form a tiny collapsed object in the core of the neutron star, under the condition of feeble repulsive DM self-interaction strength [26]. Considering the scenario related to the Chandrasekhar limit (Eq. 4 and Eq. 6), the bosonic DM particles should typically achieve the collapse criterion more easily than the fermionic ones, as mentioned in Sec. 2. However, this is not true for all the cases. The reason is, we also need to take into account the effect of $N_{\rm self}$ (see Eq. 8) which is as follows. Let us consider, a total DM mass $M_{\rm X}=N_{\rm X}m_{\chi}$ (where $N_{\rm X}$ is the total number of accumulated DM particles) is accumulated within a thermal radius of [25]
$r_{\rm th}=\left(\frac{9k\Theta}{4\pi G\rho_{\rm B}m_{\chi}}\right)^{1/2},$ (9)
@@
where $\Theta$ is the core temperature of neutron star and $\rho_{\rm B}$ is the baryon density. If the DM density [^6] Note that here, for the sake of simplicity, $N_{\rm self}$ is derived assuming that the DM density and baryon density are uniform (see Eqs. 9-11). The error introduced due to this assumption for a tiny region at the neutron star center is expected to be small. is larger than the baryon density ($\rho_{\rm B}$) within the thermal radius ($r_{\rm th}$) [25] of a host, i.e.,
$\frac{3M_{\rm X}}{4\pi r_{\rm th}^{3}}\gtrsim\rho_{\rm B},$ (10)
the DM particles become self-gravitating. Substituting Eq. (9) in Eq. (10) we obtain
$N_{\rm X}\equiv N_{\rm self}\gtrsim\frac{9}{2}\left[\frac{1}{\pi\rho_{B}m_{\chi}^{5}}\left(\frac{k\Theta}{G}\right)^{3}\right]^{1/2}$ (11)
@@
where, $k$ is the Boltzmann constant. This means that the DM becomes self-gravitating once the total number of accumulated DM particles is larger than the critical number $N_{\rm self}$. In this case, one has to consider Eqs. (4, 6, 11) carefully, and use Eq. (8) to proceed further, because the ‘Max’ function plays a crucial role here. However, for a typical neutron star, Eq. (11) reduces to
$N_{\rm self}=4.8\times 10^{41}\left(\frac{100~{}{\rm GeV}}{m_{\chi}}\right)^{5/2}\left(\frac{\Theta}{10^{5}~{}{\rm K}}\right)^{3/2}$ (12)
@@
where the temperature of the core of the neutron star is taken in the order of $10^{5}$ K [25]. We do not consider the possibility of Bose-Einstein condensate (BEC) formation here, as the temperature for BEC formation is much less [27] than the core temperature ($10^{5}$ K) of a neutron star. Now, if the captured DM particles satisfy the above-mentioned collapse criterion (Eq. 8), a tiny endoparasitic collapsed object forms at the core of the neutron star. The mass of the newly formed collapsed object should be
$M_{0}=m_{\chi}{\rm Max}[N_{\rm Ch},N_{\rm self}]<<M_{\rm h}$ (13)
@@
$t_{\rm th}\simeq 0.054~{}{\rm yr}\left(\frac{m_{\chi}}{100~{}{\rm GeV}}\right)^{2}\left(\frac{2.1\times 10^{-45}~{}{\rm cm}^{-2}}{\sigma}\right)\left(\frac{10^{5}~{}{\rm K}}{\Theta}\right)$ (15)
@@ 4 Nature of the Endoparasitic Collapsed Object
In this section, we argue that the tiny endoparasitic collapsed object is not necessarily a BH, as assumed in almost all papers (e.g., [8, 25, 27, 26]). The collapsed object could be a BH or a naked singularity depending on the density and pressure of the accumulated DM particles cloud, i.e., the initial condition of the collapsing material. As mentioned in Sec. 1, a naked singularity may arise due to an inhomogeneous collapse and non-zero pressure. It was also noted that a homogeneous collapse is an idealistic scenario. The actual process of collapse depends on many factors, such as how the captured DM particles accumulate at the stellar center after attaining the thermal equilibrium.
Let us now discuss if the DM can have pressure. It is not clear what constitute the DM in galaxy halos. There are many possibilities as shown in Fig. 1 of [24]. One belief is that the DM is pressureless (cold dark matter) [35]. However, there is no direct detection of the cold dark matter particles yet. On the other hand, the flat rotation curves observed in the outer parts of the spiral galaxies can be explained well by a halo of DM with anisotropic pressures [35]. In fact, the bending of light rays passing through the DM halo is shown to be highly sensitive to the DM pressure (see [35] and references therein). DM pressure may play a significant role in some other cases as well (see [36, 37, 39, 40, 38]). Intriguingly, it is also argued that dark energy may emerge from DM pressure [41, 42, 43] at a cosmological level. These motivate us to consider a general scenario in this paper, i.e., the pressure of the DM particles cloud can be non-zero. Furthermore, since the DM pressure could also be anisotropic [39, 38, 44, 35, 40], the collapse of the DM cloud can be anisotropic. In this context, we specifically refer to [29] (see also [30, 45]) for an interesting and detailed discussion on the pressure of bosonic matter (and the fermionic matter) and its anisotropy. Thus, the end product of the collapse of DM cloud at the center of a star can be a naked singularity.

### v2

@@ 1 Introduction
Recent gravitational wave events GW190425 [1] and GW190814 [2] have suggested the existence of near-solar-mass collapsed objects which cannot be formed via stellar evolution. These two GW events have also opened up the possibility of reliable future detections of near-solar mass collapsed objects with the advanced Laser Interferometer Gravitational-Wave Observatory (aLIGO; [3]), Lunar Gravitational-wave Antenna (LGWA; [4]) etc. An important question of physics and astronomy is how such a low mass collapsed object could form. One possibility is such near-solar mass objects are primordial black holes (PBHs; [5]), although there are observational constraints on PBHs [6]. An alternative is the capture of a PBH by a stellar host could transmute the host into a low mass black hole [7]. As the probability of such a capture is small [8], one needs to look for further alternatives. According to one such scenario, dark matter (DM) particles could accumulate around the center of a neutron star and form a mini BH by collapse, which consumes the whole star by accreting matter from it [9]. Thus, the whole neutron star is transmuted into a low mass BH.
@@
In general, a collapsed object is considered to be a BH due to the cosmic censorship conjecture (CCC) proposed by Penrose [10]. According to the CCC, the nature abhors the singularity. Otherwise, one can see the singularity from infinity, in principle, and such a collapsed object is called as a naked singularity. It was shown [11] that the initial condition of gravitational collapse determines the nature of the end product (BH or naked singularity) after completion of the gravitational collapse. For example, if the collapse is initiated by a completely homogeneous pressureless dust, with the density being constant from the center to the boundary of the collapsing cloud, the singularity will be hidden inside the event horizon, and the end product will be a BH. On the other hand, if the gravitational collapse is induced by an inhomogeneous dust, the end product can be visible as a naked singularity. Moreover, for an inhomogeneous collapsing cloud with a non-zero pressure, a naked singularity can arise [12]. Note that, as a perfectly homogeneous collapse is an idealistic scenario (e.g., stellar density changes in the radial direction), a gravitational collapse of a star may give rise to a naked singularity as the end product. In fact, Sgr A* is recently proposed as a 1st type of Joshi-Malafarina-Narayan (JMN-1 [11]) naked singularity [13]. Earlier, it was shown that M87* could be a Kerr-Taub-NUT naked singularity [14] or a Kerr naked singularity with the value of dimensionless Kerr parameter within $4.5$ to $6.5$ [14]. Similarly, GRO J1655-40 could be a Kerr-Taub-NUT naked singularity [15, 16]. In fact, it is theoretically shown in several papers [17, 12, 19, 20, 22, 21, 18] that the naked singularities are viable end products of collapse. In a recent paper [23], it is also indicated that many of the proposed primordial BHs could actually be primordial naked singularities. There are some effects already predicted [25, 24, 12] to distinguish a naked singularity from a BH, by which it may be possible to identify near- and sub-solar mass naked singularities in future. As of now, the existence of naked singularity has been hinted in a few cases as mentioned above, but there is no confirmation yet from the astrophysical observation.
@@
The nature of DM is not yet known. The interaction between the DM particles and the regular baryonic matter is via gravity, although there could be an extremely small additional interaction (with a very small cross section). There are several proposed DM candidates and some of them could be asymmetric or non-self-annihilating (see figure 1 of [26] for details). Let us consider that the non-self-annihilating DM particles enter a cosmic object/host (e.g., neutron star, white dwarf, etc.). They lose energy due to the scattering with the baryonic matter of the host. For example, in case of the neutron star, the DM particles lose energy due to the scattering with the neutron. Now, the captured DM particles are thermally distributed in a very small region of the core of the host due to the large baryon density [27]. Joining of the DM particles in an isothermal sphere around the core is known as thermalization [28]. For neutron stars, the DM reaches thermal equilibrium with the neutrons, and the captured DM particles quickly occupy a very small region near the core of the host [^1] We refer to sections III and IV of [27] for more details.. If a sufficiently large number of DM particles accumulate at the core of the host, a tiny collapsed object may form by the gravitational collapse [29]. In all the papers (e.g., [9, 29, 28, 27], and references therein) related to the dark core collapse, to the best of our knowledge, it has been presumed that the tiny collapsed object should be a BH, which may not be true. In this paper, we propose that depending on the initial condition (or, the nature) of the DM particles cloud, the tiny collapsed object could be either a BH or a naked singularity. For example, if the accumulated DM cloud is inhomogeneous in nature with a non-zero, tangential pressure, an object-like endoparasitic naked singularity [^2] Although this naked singularity is primarily hidden inside a stellar host, we call it a naked singularity in this sense that the singularity is not hidden inside the event horizon. For an object-like naked singularity, the event horizon does not form at all. could form at the core of the stellar host. The endoparasitic naked singularity should start accreting matter from its host immediately after its formation, and transmute the entire host into a naked singularity. Note that it is recently proposed that the gravitational collapse also gives rise to the primordial naked singularities [23] instead of primordial BHs depending on the nature of the collapsing matter. If such a primordial naked singularity is captured by a stellar object, it may transmute into a collapsed object. However, here we discuss only the transmutation led by an endoparasitic naked singularity formed by the accumulation of DM particles.
@@
This paper is organized as follows. We discuss the Chandrasekhar limit in section 2. section 3 describes the formation of an endoparasitic collapsed object inside a neutron star. We show in section 4 that the endoparasitic collapsed object could be a BH or a naked singularity depending on the initial conditions of the DM cloud. In section 5 we discuss the transmutation timescale, and we argue on the nature (BH or naked singularity) of the transmuted collapsed object in section 6. We constrain a parameter space using the ‘missing pulsar’ problem in section 7, and conclude in section 8.
@@ 2 Chandrasekhar limit
$E_{\rm F}\sim\hbar cn^{1/3}\sim\hbar c\frac{N^{1/3}}{R}$ (2.1)
@@
$E=E_{\rm F}+E_{\rm G}=\hbar c\frac{N^{1/3}}{R}-\frac{GNm^{2}}{R}.$ (2.3)
@@
$N_{\rm Ch}^{\rm f}\sim\left(\frac{\hbar c}{Gm^{2}}\right)^{3/2}=\left(\frac{M_{\rm P}}{m}\right)^{3}=1.8\times 10^{51}\left(\frac{100~{}{\rm GeV}}{m}\right)^{3},$ (2.4)
@@
$E=\hbar c\frac{1}{R}-\frac{GNm^{2}}{R}.$ (2.5)
@@
$N_{\rm Ch}^{\rm b}\sim\left(\frac{\hbar c}{Gm^{2}}\right)=\left(\frac{M_{\rm P}}{m}\right)^{2}=1.5\times 10^{34}\left(\frac{100~{}{\rm GeV}}{m}\right)^{2}.$ (2.6)
For both fermions and bosons, the gravitational collapse eventually leads to form a collapsed object. Comparing eqs. (2.4) and (2.6), one can see that bosons usually collapse more easily than fermions.
@@ 3 Formation of Endoparasitic Collapsed Objects inside Neutron stars
$C=\sqrt{\frac{6}{\pi}}\frac{\rho_{\rm DM}}{0.45m_{\chi}}{2GM_{\rm h}R_{\rm h}\over\bar{v}_{\chi}}.\left[1-\exp{\left(-\frac{\sigma}{\sigma_{\rm cr}}\right)}\right]\left[1-\frac{1-\exp{(-\xi)}}{\xi}\right]$ (3.1)
@@
where $N$ is the total number of DM particles accumulated inside the neutron star. $N_{\rm Ch}$ (see eqs. 2.4 and 2.6) and $N_{\rm self}$ are the Chandrasekhar limit and the number of DM particles required for initiating the self-gravitating collapse, respectively [27]. If the total number ($N$) of fermions(f)/bosons(b) increases beyond $N^{\rm f}_{\rm Ch}$ or $N^{\rm b}_{\rm Ch}$ [27], the gravitational energy dominates over the total particle energy, and, the gravitational collapse occurs to form a tiny collapsed object in the core of the neutron star, under the condition of feeble repulsive DM self-interaction strength [28]. Considering the scenario related to the Chandrasekhar limit (eq. 2.4 and eq. 2.6), the bosonic DM particles should typically achieve the collapse criterion more easily than the fermionic ones, as mentioned in section 2. However, this is not true for all the cases. The reason is, we also need to take into account the effect of $N_{\rm self}$ (see eq. 3.2) which is as follows. Let us consider, a total DM mass $M_{\rm X}=N_{\rm X}m_{\chi}$ (where $N_{\rm X}$ is the total number of accumulated DM particles) is accumulated within a thermal radius of [27]
$r_{\rm th}=\left(\frac{9k\Theta}{4\pi G\rho_{\rm B}m_{\chi}}\right)^{1/2},$ (3.3)
@@
where $\Theta$ is the core temperature of neutron star and $\rho_{\rm B}$ is the baryon density. If the DM density [^3] Note that here, for the sake of simplicity, $N_{\rm self}$ is derived assuming that the DM density and baryon density are uniform (see eqs. 3.3-3.5). The error introduced due to this assumption for a tiny region at the neutron star center is expected to be small. is larger than the baryon density ($\rho_{\rm B}$) within the thermal radius ($r_{\rm th}$) [27] of a host, i.e.,
$\frac{3M_{\rm X}}{4\pi r_{\rm th}^{3}}\gtrsim\rho_{\rm B},$ (3.4)
the DM particles become self-gravitating. Substituting eq. (3.3) in eq. (3.4) we obtain
$N_{\rm X}\equiv N_{\rm self}\gtrsim\frac{9}{2}\left[\frac{1}{\pi\rho_{B}m_{\chi}^{5}}\left(\frac{k\Theta}{G}\right)^{3}\right]^{1/2}$ (3.5)
@@
where, $k$ is the Boltzmann constant. This means that the DM becomes self-gravitating once the total number of accumulated DM particles is larger than the critical number $N_{\rm self}$. In this case, one has to consider eqs. (2.4, 2.6, 3.5) carefully, and use eq. (3.2) to proceed further, because the ‘Max’ function plays a crucial role here. However, for a typical neutron star, eq. (3.5) reduces to
$N_{\rm self}=4.8\times 10^{41}\left(\frac{100~{}{\rm GeV}}{m_{\chi}}\right)^{5/2}\left(\frac{\Theta}{10^{5}~{}{\rm K}}\right)^{3/2}$ (3.6)
@@
where the temperature of the core of the neutron star is taken in the order of $10^{5}$ K [27]. We do not consider the possibility of Bose-Einstein condensate (BEC) formation here, as the temperature for BEC formation is much less [29] than the core temperature ($10^{5}$ K) of a neutron star. Now, if the captured DM particles satisfy the above-mentioned collapse criterion (eq. 3.2), a tiny endoparasitic collapsed object forms at the core of the neutron star. The mass of the newly formed collapsed object should be
$M_{0}=m_{\chi}{\rm Max}[N_{\rm Ch},N_{\rm self}]<<M_{\rm h}$ (3.7)
@@
$t_{\rm th}\simeq 0.054~{}{\rm yr}\left(\frac{m_{\chi}}{100~{}{\rm GeV}}\right)^{2}\left(\frac{2.1\times 10^{-45}~{}{\rm cm}^{-2}}{\sigma}\right)\left(\frac{10^{5}~{}{\rm K}}{\Theta}\right)$ (3.9)
@@ 4 Nature of the Endoparasitic Collapsed Object
In this section, we argue that the tiny endoparasitic collapsed object is not necessarily a BH, as assumed in almost all papers (e.g., [9, 27, 29, 28]). The collapsed object could be a BH or a naked singularity depending on the density and pressure of the accumulated DM particles cloud, i.e., the initial condition of the collapsing material. As mentioned in section 1, a naked singularity may arise due to an inhomogeneous collapse and non-zero pressure. It was also noted that a homogeneous collapse is an idealistic scenario. The actual process of collapse depends on many factors, such as how the captured DM particles accumulate at the stellar center after attaining the thermal equilibrium.
Let us now discuss if the DM can have pressure. It is not clear what constitute the DM in galaxy halos. There are many possibilities as shown in figure 1 of [26]. One belief is that the DM is pressureless (cold dark matter) [37]. However, there is no direct detection of the cold dark matter particles yet. On the other hand, the flat rotation curves observed in the outer parts of the spiral galaxies can be explained well by a halo of DM with anisotropic pressures [37]. In fact, the bending of light rays passing through the DM halo is shown to be highly sensitive to the DM pressure (see [37] and references therein). DM pressure may play a significant role in some other cases as well (see [38, 39, 41, 42, 40]). Intriguingly, it is also argued that dark energy may emerge from DM pressure [43, 44, 45] at a cosmological level. These motivate us to consider a general scenario in this paper, i.e., the pressure of the DM particles cloud can be non-zero. Furthermore, since the DM pressure could also be anisotropic [41, 40, 46, 37, 42], the collapse of the DM cloud can be anisotropic. In this context, we specifically refer to [31] (see also [32, 47]) for an interesting and detailed discussion on the pressure of bosonic matter (and the fermionic matter) and its anisotropy. Thus, the end product of the collapse of DM cloud at the center of a star can be a naked singularity.